Help...
A sailboat travels 20 miles downstream in 4 hours. It returns in 5 hours. Find the speed of the sailboat in still water and the rate of the current.
Ugggg
Thanks

Answer by Alan3354(69443)

A sailboat travels 20 miles downstream in 4 hours. It returns in 5 hours. Find the speed of the sailboat in still water and the rate of the current
--------------------
Write down what you know:
b = speed of the boat
c = speed of the current
Downstream the speed is b+c
Upstream it's b-c
The distance is the same
---------------------------
d = rt
20 = 4*(b+c) = 5*(b-c)
-----------------------
4b + 4c = 20
5b - 5c = 20
------------
b + c = 5
b - c = 4
------------ Add
2b = 9
b = 4.5 mph
------------
c = 1/2 mph